# Universality of Estimator for High-Dimensional Linear Models with Block Dependency

### v1

@@ 3.1. Gaussian Analogy
Note that numerous studies has investigated the asymptotic behaviour of the estimator and the risk in the proportionally high-dimensional limit.
@@ 3.2. Universality of Empirical Risk
$\left|E\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{X})\right)\right]-E\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{G})\right)\right]\right|\leq C_{0}\cdot K_{g}\cdot r_{R}(L_{n}).$
@@
$r_{R}(L_{n}):=\inf_{\rho\in(0,\bar{\rho})}\left\{\mathcal{M}_{\ell}(\rho)+\mathcal{D}^{3}_{\ell}(\rho)\inf_{\delta\in(0,\omega_{n})}\left[\mathcal{M}_{R}(L_{n},\delta)+Av(\{L_{\ell_{i}}^{\mathbf{q}}\})^{1/3}\cdot L_{n}^{(\bar{\mathbf{q}}_{0}+3)/{3}}\log^{2/3}_{+}(L_{n}/\delta)\sigma^{1/3}_{n}\right]\right\}$
@@
Here $C_{1}>0$ is an absolute multiple of $C_{0}$.
Theorem 1 shows that the difference in expectations of the minimized risk functions is characterized by $r_{R}(L_{n})$, which includes the properties of the loss function $\ell_{i}$ and the regularizer $R$. The asymptotic equivalence of the distributions is guaranteed if $r_{R}(L_{n})$ converges to to zero, which largely on the convergence of $\sigma_{n}$.
From the result of Theorem 1, we can discuss the value of $d$ required to converge the error to zero. The following corollary summarizes the result:
@@
Consider one of the loss functions presented in Example 1. Suppose that the conditions in Theorem 1, $\bar{\mathbf{q}}_{0}\leq 6$, and $Av(\{L_{\ell_{i}}^{\mathbf{q}}\})=O(1)$ hold. Set $L_{n}=O(n^{\alpha}\log^{\kappa}n)$ with $\alpha\in[0,1/(2\bar{\mathbf{q}}_{0}+6)]$ and $\kappa\geq 0$ as $n\to\infty$. Then, with setting $d=o(n^{1/4-(\bar{\mathbf{q}}_{0}+3)\alpha/2}/\log^{6+(2\bar{\mathbf{q}}_{0}+6)\kappa}n)$ as $n\to\infty$, for any $\mathcal{S}_{n}\subset[-L_{n},L_{n}]^{p}$ with $L_{n}\geq 1$, and any $g\in C^{3}(\mathbb{R})$, we have the following as $n,p\to\infty$:
@@
We first note the setup of $L_{n}$. As seen from Section 3 in Han and Shen (2023) or Section 7 in Lahiry and Sur (2024), typical estimator examples satisfy $\bar{\mathbf{q}}_{0}\leq 6$, and also $L_{n}=O(\log^{\kappa}n)$ with $\kappa\geq 0$ is enough to cover a support of distributions of the estimators. In this situation, we can set $\alpha=0$ and obtain an order $d=o(n^{1/4})$ up to logarithmic factors that makes the discrepancy converge to zero in the proportionally high-dimensional limit.
We discuss the relationship between Theorem 1 and previous results of the existing studies. Theorem 1 can be regarded as a generalization of Theorem 2.3 in Han and Shen (2023). Han and Shen (2023) considers the case where all elements of $X_{i}$ are independent, i.e., $d$-block dependent with $d=1$, and the statement of their theorem is equivalent to Theorem 1 with $\sigma_{n}$ replaced by $n^{-1/2}$. Additionally, Theorem 4.1 of Lahiry and Sur (2024) has a similar universality statement under the block dependence, where $d=o((n/\log n)^{1/5})$ holds. Note that the permissible order of Lahiry and Sur (2024) is more restrictive than those allowed in Theorem 1 and Corollary 2 with $\alpha=0$. Furthermore, Chang et al. (2024) shows the universality of a maximum of high-dimensional random vector with the dependent random variables. Although there is some technical similarity, the result by Chang et al. (2024) is not comparable with our result, because their setup is different from our setting. For details, see Remark 2.
@@ 3.3. Universality of Estimator
    $\max\left\{\mathbb{P}(\|\widehat{w}_{\mathbf{G}}\|_{\infty}>L_{n}),\mathbb{P}(\|\widehat{w}_{\mathbf{X}}\|_{\infty}>L_{n})\right\}\leq\varepsilon_{n}$
@@
$\mathbb{P}\left(\widehat{w}_{\mathbf{X}}\in\mathcal{S}_{n}\right)\leq 4\varepsilon_{n}+C_{0}\max\{1,\rho_{0}^{-3}\}r_{R}(L_{n})$
@@
Consider one of the loss functions presented in Example 1. Suppose that the conditions of Theorem 3 hold with a fixed $\mathcal{S}_{n}\subset\mathbb{R}^{p}$ and $\varepsilon_{n}\in[0,1/4)$. $\bar{\mathbf{q}}_{0}\leq 6$, and $Av(\{L_{\ell_{i}}^{\mathbf{q}}\})=O(1)$ hold. Set $L_{n}=O(n^{\alpha}\log^{\kappa}n)$ with $\alpha\in[0,1/(2\bar{\mathbf{q}}_{0}+6)]$ and $\kappa\geq 0$ as $n\to\infty$. Then, with setting $d=o(n^{1/4-(\bar{\mathbf{q}}_{0}+3)\alpha/2}/\log^{6+(2\bar{\mathbf{q}}_{0}+6)\kappa}n)$ as $n\to\infty$, we have the following as $n,p\to\infty$:
@@
We provide explanations on conditions (i) and (ii) in this Theorem 3. Condition (i) requires the boundedness of the estimators $\widehat{w}_{\mathbf{X}}$ and $\widehat{w}_{\mathbf{G}}$, which is demonstrated depending on each example. Condition (ii) can be investigated by a limiting value of the minimum of the empirical risk in the proportionally high-dimension regime with Gaussian covariate, which also depends on specific examples. These two conditions are identical to those provided in Theorem 2.4 of Han and Shen (2023), and more detailed discussions will be given in the subsequent section on applications.
@@ 4. Outline of Proof with Lindeberg Principle
In this section, we outline the proof of Theorem 1. To emphasize the fact that Theorem 1 still holds even though the Gaussianity of $\mathbf{G}$ is violated, we replace $\mathbf{G}$ with another design matrix $\mathbf{W}=(W_{1},...,W_{n})^{\top}\in\mathbb{R}^{n\times p}$ that has the same first and second moments as $\mathbf{X}$, that is, $\mathbb{E}[W_{i}]=\mathbb{E}[X_{i}]=0$ and $\mathbb{E}[W_{i}W_{i}^{\top}]=\mathbb{E}[X_{i}X_{i}^{\top}]=\Sigma/n$ hold for $i=1,...,n$.
@@ 4.2.1. Generalized Lindeberg Principle
We develop a generalized Lindberg principle, which is a method used in evaluating discrepancy between random vectors by swapping their elements and evaluating the error using the Taylor approximation. Specifically, we employ this principle to evaluate the discrepancy of random matrices with block dependence. The independence of $B_{j}$ from any other $B_{i}$ allows us to upper bound the difference in expectations for any function $f\in C^{3}(\mathbb{R}^{n\times p})$.
We provide a key result of the Lindeberg principle with the block dependence. We define $p$-dimensional random vector $X_{i}^{(B_{k})}:=(X_{i,1}\textbf{1}\{1\in B_{k}\},...,X_{i,p}\textbf{1}\{p\in B_{k}\})^{\top}$. Then, we obtain the following lemmas.
@@
Consider two sequences of i.i.d. $\mathbb{R}^{p}$-valued random vectors $X_{1},...,X_{n}$ and $W_{1},...,W_{n}$. Suppose that each $\{X_{i}\}_{i=1}^{n}$ and $\{W_{i}\}_{i=1}^{n}$ is $d$-block dependent, and $E[X_{i}]=E[W_{i}]$ and $E[X_{i}X_{i}^{\top}]=E[W_{i}W_{i}^{\top}]$ hold. Then, for any $f\in C^{3}(\mathbb{R}^{n\times p})$ satisfying $\max_{i\in[n],j\in[p]}|E[\partial_{i,j}f(\mathbf{Z}^{0}_{i,j})]|<\infty$ and $\max_{i\in[n],j,k\in[p]}|E[\partial_{i,j}\partial_{i,k}f(\mathbf{Z}^{0}_{i,j})]|<\infty$, we have
@@
This result indicates that the difference in the distributions of $\mathbf{X}$ and $\mathbf{W}$ is evaluated based on the size of $B_{j}$ of the block dependencies. Since there is a triple summation over $B_{j}$ within the upper bound of the lemma, the cubic value $|B_{j}|^{3}$ becomes critical. If we have $|B_{j}|=1$, which corresponds to the case where all elements of $X_{i}$ and $W_{i}$ are independent, the expectation terms of the upper bound remains bounded. In the following sections, it becomes crucial to assess $\sum_{j}|B_{j}|^{m}$ with $m\geq 0$ so that this upper bound does not diverge.
@@ 4.2.2. Bounding Sum of Block Size
In the case of $m=1$, $\sum_{j=1}^{k}|B_{j}|=p$ holds by its definition. Hence, we obtain the statement.
In the case of $m\geq 2$, to obtain the statement, we generate newly aligned cells $D_{1},...,D_{s}\subset[p]$ from the original cells of the partition $\{B_{j}\}_{j=1}^{k}$ by the following steps described by a pseudo code:
  1. 1.
    Define $I:=\{\ell\in[k]\mid|B_{\ell}|\geq\lfloor d/2\rfloor+1\}$ and $I^{c}:=[k]\backslash I$.
  2. 2.
    Initialize $i\leftarrow 1$.
  3. 3.
    For $j\in I$, repeat the followings:
    1. i.
      Define $D_{i}:=B_{j}$.
    2. ii.
      Set $i\leftarrow i+1$.
  4. 4.
    For $j\in I^{c}$, repeat the followings:
    1. i.
      Define $D_{i}:=B_{j}$.
    2. ii.
      Set $I^{c}\leftarrow I^{c}\backslash\{j\}$.
    3. iii.
      If $I^{c}=\emptyset$, then break.
    4. iv.
      For $j^{\prime}\in I^{c}$, repeat the followings:
      1. a.
        Set $I^{c}\leftarrow I^{c}\backslash\{j^{\prime}\}$.
      2. b.
        Set $D_{i}\leftarrow D_{i}\cup B_{j^{\prime}}$.
      3. c.
        If $|D_{i}|\geq\lfloor d/2\rfloor+1$ or $I^{c}=\emptyset$, then break.
    5. v.
      If $I^{c}=\emptyset$, then break:
    6. vi.
      Set $i\leftarrow i+1$.
  5. 5.
    Define $s\leftarrow i$.
  6. 6.
    Output $D_{1},D_{2},...,D_{s}$.
From the steps, the generated cells $D_{1},D_{2},\cdots,D_{s}\subset[p]$ satisfy $\sum_{j=1}^{k}|B_{j}|^{m}\leq\sum_{\ell=1}^{s}|D_{\ell}|^{m}$. Furthermore, it is clear from the algorithm that at most one cell $D_{j}$ satisfies $|D_{j}|\leq\lfloor d/2\rfloor$, while all the other cells satisfy $|D_{j}|\geq\lfloor d/2\rfloor+1$. Hence, $s$ satisfies the following inequality:
$s$ $\leq 1+\lceil p/(\lfloor d/2\rfloor+1)\rceil\leq 1+\lceil 2p/d\rceil\leq 2p/d+2\leq 4(p/d),$
where the last inequality holds from the assumptions that $d\leq p$. Therefore, we have
$\sum_{j=1}^{k}|B_{j}|^{m}$ $\leq\sum_{\ell=1}^{s}|D_{\ell}|^{m}\leq\sum_{\ell=1}^{s}d^{m}\leq 4pd^{m-1}.$
The important consequence in Lemma 6 is that the upper bound holds independently of $k$. In other words, as long as we determine the upper bound $d$ of the block size, we do not need to care how many blocks are used to divide the $p$ variables. In the proof, we arrange the existing cells to construct new cells, which are then used to control the total sum.
@@ 4.2.3. Unifying These Results
Using Lemma 5 and Lemma 6 above, we bound the term $IV$, which describes the difference between the distributions of the minimized risk with the datasets. The statement follows $g\in C^{3}(\mathbb{R}^{n\times p})$.
@@ 4.3.2. Bound for $II$: Discretization of Parameter Space
We study the term $II$, which explains an effect of approximating the minimization problem $\min_{w\in\mathcal{S}_{n}}H_{\ell{\rho}}(w,\mathbf{X})$ on the parameter space $\mathcal{S}_{n}$ to the minimization problem $\min_{w\in\mathcal{S}_{n,\delta}}H_{\ell{\rho}}(w,\mathbf{X})$ on a discretized parameter space $\mathcal{S}_{n,\delta}$. To the aim, we obtain the following statement.
@@
$\leq C_{0}\|g^{\prime}\|_{\infty}\left(\mathcal{D}_{\ell}(\rho)\delta\cdot L_{n}^{q_{0}q_{1}}Av({L_{\ell_{i}}^{q_{1}}})\cdot\left(E\left[\max_{1\leq i\leq n}\sum_{j=1}^{p}|X_{i,j}|\right]+E\left[\left(\max_{1\leq i\leq n}\sum_{j=1}^{p}|X_{i,j}|\right)^{q_{0}q_{1}+1}\right]\right)+\mathcal{M}_{R}(L_{n},\delta)\right).$

### v2

@@ 3.1. Gaussian analogy
Note that numerous studies have investigated the asymptotic behavior of the estimator and the risk in the proportionally high-dimensional limit.
@@ 3.2. Universality of empirical risk
$\left|E\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{X})\right)\right]-E\left[g\left(\min_{w\in\mathcal{S}_{n}}H_{\ell}(w,\mathbf{G})\right)\right]\right|\leq C_{0}\cdot K_{g}\cdot r_{R}(L_{n}).$ (7)
@@
$r_{R}(L_{n}):=\inf_{\rho\in(0,\bar{\rho})}\Biggl\{$ $\mathcal{M}_{\ell}(\rho)+\mathcal{D}^{3}_{\ell}(\rho)$
$\times\inf_{\delta\in(0,\omega_{n})}\left[\mathcal{M}_{R}(L_{n},\delta)+Av(\{L_{\ell_{i}}^{\mathbf{q}}\})^{1/3}\cdot L_{n}^{(\bar{\mathbf{q}}_{0}+3)/{3}}\log^{2/3}_{+}(L_{n}/\delta)\sigma^{1/3}_{n}\right]\Biggr\},$
@@
Here $C_{1}>0$ is a constant multiple of $C_{0}$.
Theorem 1 shows that the difference in expectations of the minimized risk functions is characterized by $r_{R}(L_{n})$, which includes the properties of the loss function $\ell_{i}$ and the regularizer $R(\cdot)$. The asymptotic equivalence of the distributions is guaranteed if $r_{R}(L_{n})$ converges to zero, which largely depends on the convergence of $\sigma_{n}$.
Note that the inequality (7) is conditional on the noise variable $\xi_{1},...,\xi_{n}$: the left-hand side is a conditional expectation on the noise and the term $r_{R}(L_{n})$ in the right-hand side is a $\xi_{i}$-dependent random variable. Even in this case, we can give a non-probabilistic upper bound on this term using the weak law of large numbers reflecting the shape of the loss function. The specific upper bound is given in Section 5 for the applications.
From the result of Theorem 1, we can discuss the value of $d$ required for the error to converge to zero. The following corollary summarizes the result:
@@
Consider one of the loss functions presented in Example 1 and the regularizer $R(\cdot)=\|\cdot\|_{1}$ or $R(\cdot)=\|\cdot\|_{2}^{2}$. Suppose that the conditions in Theorem 1, $\bar{\mathbf{q}}_{0}\leq 6$, and $Av(\{L_{\ell_{i}}^{\mathbf{q}}\})=O(1)$ hold. Set $L_{n}=O(n^{\alpha}\log^{\kappa}n)$ with $\alpha\in[0,1/(2\bar{\mathbf{q}}_{0}+6)]$ and $\kappa\geq 0$ as $n\to\infty$. Then, with setting $d=o(n^{1/4-(\bar{\mathbf{q}}_{0}+3)\alpha/2}/\log^{3/2+1/2(\bar{\mathbf{q}}_{0}+3)\kappa}n)$ as $n\to\infty$, for any $\mathcal{S}_{n}\subset[-L_{n},L_{n}]^{p}$ with $L_{n}\geq 1$, and any $g\in C^{3}(\mathbb{R})$, we have the following as $n,p\to\infty$:
@@
We first note the setup of $L_{n}$. As seen from Section 3 in Han and Shen (2023) or Section 7 in Lahiry and Sur (2024), typical estimator examples satisfy $\bar{\mathbf{q}}_{0}\leq 6$, and also $L_{n}=O(\log^{\kappa}n)$ with $\kappa\geq 0$ is enough to cover the support of the estimator distribution. In this situation, we can set $\alpha=0$ and obtain an order $d=o(n^{1/4})$ up to logarithmic factors that makes the discrepancy converge to zero in the proportionally high-dimensional limit.
We discuss the relationship between Theorem 1 and previous results of the existing studies. Theorem 1 can be regarded as a generalization of Theorem 2.3 in Han and Shen (2023). Han and Shen (2023) considers the case where all elements of $X_{i}$ are independent, i.e., $d$-block dependent with $d=1$, and the statement of their theorem is equivalent to Theorem 1 with $\sigma_{n}$ replaced by $n^{-1/2}$. Additionally, Theorem 4.1 of Lahiry and Sur (2024) has a similar universality statement under the block dependence, where $d=o((n/\log n)^{1/5})$ holds. Note that the permissible order of Lahiry and Sur (2024) is more restrictive than that allowed in Theorem 1 and Corollary 2 with $\alpha=0$. Furthermore, Chang et al. (2024) shows the universality of the maximum of high-dimensional random vectors with dependent random variables. Although there is some technical similarity, the result by Chang et al. (2024) is not comparable with our result, because their setup is different from our setting. For details, see Remark 2.
@@ 3.3. Universality of estimator
    $\max\left\{\mathbb{P}(\|\widehat{w}_{\mathbf{G}}\|_{\infty}>L_{n}),\mathbb{P}(\|\widehat{w}_{\mathbf{X}}\|_{\infty}>L_{n})\right\}\leq\varepsilon_{n},$
@@
$\mathbb{P}\left(\widehat{w}_{\mathbf{X}}\in\mathcal{S}_{n}\right)\leq 4\varepsilon_{n}+C_{0}\max\{1,\rho_{0}^{-3}\}r_{R}(L_{n}).$
@@
Consider one of the loss functions presented in Example 1 and the regularizer $R(\cdot)=\|\cdot\|_{1}$ or $R(\cdot)=\|\cdot\|_{2}^{2}$. Suppose that the conditions of Theorem 3 hold with a fixed $\mathcal{S}_{n}\subset\mathbb{R}^{p}$ and $\varepsilon_{n}\in[0,1/4)$. $\bar{\mathbf{q}}_{0}\leq 6$, and $Av(\{L_{\ell_{i}}^{\mathbf{q}}\})=O(1)$ hold. Set $L_{n}=O(n^{\alpha}\log^{\kappa}n)$ with $\alpha\in[0,1/(2\bar{\mathbf{q}}_{0}+6)]$ and $\kappa\geq 0$ as $n\to\infty$. Then, with setting $d=o(n^{1/4-(\bar{\mathbf{q}}_{0}+3)\alpha/2}/\log^{3/2+1/2(\bar{\mathbf{q}}_{0}+3)\kappa}n)$ as $n\to\infty$, we have the following as $n,p\to\infty$:
@@
We provide explanations on conditions (i) and (ii) in Theorem 3. Condition (i) requires the boundedness of the estimators $\widehat{w}_{\mathbf{X}}$ and $\widehat{w}_{\mathbf{G}}$, which is demonstrated depending on each example. Condition (ii) can be investigated by a limiting value of the minimum of the empirical risk in the proportionally high-dimension regime with a Gaussian covariate, which also depends on specific examples. These two conditions are identical to those provided in Theorem 2.4 of Han and Shen (2023), and more detailed discussions will be given in the subsequent section on applications.
@@ 4. Outline of proof with Lindeberg principle
In this section, we outline the proof of Theorem 1. To emphasize the fact that Theorem 1 still holds even though the Gaussianity of $\mathbf{G}$ is violated, we replace $\mathbf{G}$ with another design matrix $\mathbf{W}=(W_{1},...,W_{n})^{\top}\in\mathbb{R}^{n\times p}$ that has the same first and second moments as $\mathbf{X}$, that is, $\mathbb{E}[W_{i}]=\mathbb{E}[X_{i}]=0$ and $\mathbb{E}[W_{i}W_{i}^{\top}]=\mathbb{E}[X_{i}X_{i}^{\top}]=\Sigma/n$ hold for i=1,,ni=1,...,n.
@@ 4.2.1. Generalized Lindeberg principle
We develop a generalized Lindeberg principle, which is a method used in evaluating a discrepancy between random vectors by swapping their elements and evaluating the error using the Taylor approximation. Specifically, we employ this principle to evaluate the discrepancy of random matrices with block dependence. The independence of $B_{j}$ from any other $B_{i}$ allows us to upper bound the difference in expectations for any function $f\in C^{3}(\mathbb{R}^{n\times p})$.
We provide a key result of the Lindeberg principle with the block dependence. We define $p$-dimensional random vector $X_{i}^{(B_{k})}:=(X_{i,1}\textbf{1}\{1\in B_{k}\},...,X_{i,p}\textbf{1}\{p\in B_{k}\})^{\top}\in\mathbb{R}^{p}$. Then, we obtain the following lemmas.
@@
Consider two sequences of i.i.d. $\mathbb{R}^{p}$-valued random vectors $X_{1},...,X_{n}$ and $W_{1},...,W_{n}$. Suppose that each $\{X_{i}\}_{i=1}^{n}$ and $\{W_{i}\}_{i=1}^{n}$ is $d$-block dependent, and both $E[X_{i}]=E[W_{i}]$ and $E[X_{i}X_{i}^{\top}]=E[W_{i}W_{i}^{\top}]$ hold. Then, for any $1\leq j\leq k$ and $f\in C^{3}(\mathbb{R}^{n\times p})$ satisfying both inequalities $\max_{i\in[n],m\in[p]}|E[\partial_{i,m}f(\mathbf{Z}^{0}_{i,j})]|<\infty$ and $\max_{i\in[n],m,m^{\prime}\in[p]}|E[\partial_{i,m}\partial_{i,m^{\prime}}f(\mathbf{Z}^{0}_{i,j})]|<\infty$, we have
@@
This lemma is technically important for our result. Specifically, we bound the difference $|E[f(\mathbf{X})]-E[f(\mathbf{W})]|$ by the Taylor expansion, then we can cancel out interactions between variables belonging to different cells $B_{j}$ using the block dependence. Thus, if the cells $B_{j}$ are not excessively large, we can properly control for the effects of the dependencies.
This result indicates that the difference in the distributions of $\mathbf{X}$ and $\mathbf{W}$ is evaluated based on the size of $B_{j}$ of the block dependencies. Since there is a triple summation over $B_{j}$ within the upper bound of the lemma, the cubic value $|B_{j}|^{3}$ becomes critical. If we have $|B_{j}|=1$, which corresponds to the case where all elements of $X_{i}$ and $W_{i}$ are independent, the expectation terms of the upper bound remain bounded. In the following sections, it becomes crucial to assess $\sum_{j}|B_{j}|^{m}$ with $m\geq 0$ so that this upper bound does not diverge.
@@ 4.2.2. Bounding sum of block size
A straightforward calculation gives
$\sum_{j=1}^{k}|B_{j}|^{m}$ $=\sum_{j=1}^{k}|B_{j}||B_{j}|^{m-1}\leq\sum_{j=1}^{k}|B_{j}|d^{m-1}=pd^{m-1}.$
where the inequality holds from the assumption $\max_{j=1,...,k}|B_{j}|\leq d$ and the last equality follows from the property of cells. ∎
The important consequence in Lemma 6 is that the upper bound holds independently of $k$. In other words, as long as we determine the upper bound $d$ of the block size, we do not need to care how many blocks are used to divide the $p$ variables. The equality holds when $k=p/d$ with $|B_{j}|=d$ for all j=1,,kj=1,...,k.
@@ 4.2.3. Unifying these results
Using Lemma 5 and Lemma 6 above, we bound the term $IV$, which describes the difference between the distributions of the minimized risk with the datasets. The statement holds for $g\in C^{3}(\mathbb{R}^{n\times p})$.
@@ 4.3.2. Bound for $II$: discretization of parameter space
We study the term $II$, which explains an effect of approximating the minimization problem $\min_{w\in\mathcal{S}_{n}}H_{\ell{\rho}}(w,\mathbf{X})$ on the parameter space $\mathcal{S}_{n}$ to the minimization problem $\min_{w\in\mathcal{S}_{n,\delta}}H_{\ell{\rho}}(w,\mathbf{X})$ on a discretized parameter space $\mathcal{S}_{n,\delta}$. To this aim, we obtain the following statement.
@@
$\leq C_{0}\|g^{\prime}\|_{\infty}\left(\mathcal{D}_{\ell}(\rho)\delta\cdot L_{n}^{q_{0}q_{1}}Av({L_{\ell_{i}}^{q_{1}}})\cdot\mathcal{M}_{X}+\mathcal{M}_{R}(L_{n},\delta)\right),$
where $\mathcal{M}_{X}:=E[\max_{1\leq i\leq n}\sum_{j=1}^{p}|X_{i,j}|]+E[(\max_{1\leq i\leq n}\sum_{j=1}^{p}|X_{i,j}|)^{q_{0}q_{1}+1}]$.